OpenStudy (anonymous):

logx+log(x/10)+log(1/19)=1 all are to base 9 solve

OpenStudy (anonymous):

x=sqrt190

OpenStudy (anonymous):

thanks for the DETAILED explanation

OpenStudy (anonymous):

Since log a + log b = log (a*b) and log a to base a = 1

OpenStudy (anonymous):

sorry sqrt 190 is wrong

OpenStudy (anonymous):

correct one is sqrt190*9

OpenStudy (anonymous):

sorry for not giving detailed solution.

OpenStudy (anonymous):

method: log ( x * x/10 * 1/19)= log 9. x^2/190=9 x=sqrt(190*9)

OpenStudy (anonymous):

Sorry for miscalculation earlier. I wad half asleep

OpenStudy (anonymous):

*was

OpenStudy (anonymous):

to solve the eqn we have to eliminate log. u can write 1=log9(base is 9) and use loga + logb=loga*b.on simplifying u get logxsquared/109=log9 and u get the answer